# DynamicVL: Benchmarking Multimodal Large Language Models for Dynamic City Understanding

### v1

@@ Abstract
Multimodal large language models have demonstrated remarkable capabilities in visual understanding, but their application to long-term Earth observation analysis remains limited, primarily focusing on single-temporal or bi-temporal imagery. To address this gap, we introduce DVL-Suite, a comprehensive framework for analyzing long-term urban dynamics through remote sensing imagery. Our suite comprises 15,063 high-resolution (1.0m) multi-temporal images spanning 42 megacities in the U.S. from 2005 to 2023, organized into two components: DVL-Bench and DVL-Instruct. The DVL-Bench includes seven urban understanding tasks, from fundamental change detection (pixel-level) to quantitative analyses (regional-level) and comprehensive urban narratives (scene-level), capturing diverse urban dynamics including expansion/transformation patterns, disaster assessment, and environmental challenges. We evaluate 17 state-of-the-art multimodal large language models and reveal their limitations in long-term temporal understanding and quantitative analysis. These challenges motivate the creation of DVL-Instruct, a specialized instruction-tuning dataset designed to enhance models’ capabilities in multi-temporal Earth observation. Building upon this dataset, we develop DVLChat, a baseline model capable of both image-level question-answering and pixel-level segmentation, facilitating a comprehensive understanding of city dynamics through language interactions.
@@ 1 Introduction
Sustainable city, as a key goal in “The 2030 Agenda for Sustainable Development” [^1] https://sdgs.un.org/goals/goal11, has proposed new requirements for urban resilience, convenience, and comfort. Remote sensing technology enables us to monitor urban development over time by analyzing satellite imagery, allowing us to track large-scale changes in urban landscapes frolking2024global; xiong2024earthnets. However, research in this field has been largely limited to comparing images from just two time points zhengsegment; wu2023fully, primarily due to the scarcity of well-aligned vision-language datasets spanning longer time series. This limitation has constrained our ability to conduct a comprehensive, large-scale understanding of urban dynamics.
@@
The recent emergence of Multimodal Large Language Models (MLLMs) li2024llava; liu2024llavanext; wang2024qwen2vl represents a significant advancement in visual-language understanding. These models mark a shift from specialized, single-purpose systems to versatile frameworks capable of handling multiple tasks, including but not limited to visual grounding liu2024bench, image captioning dai2023instructblip, and visual question answering yue2024mmmu; yue2024mmmupro. While recent MLLM research has demonstrated promising results in multi-image meng2024mmiu; jiangmantis and video understanding wu2024longvideobench; hu2025videommmu, these efforts primarily focus on daily scenes. In the context of remote sensing, existing approaches irvin2024teochat; soni2024earthdial to multi-temporal analysis are typically limited to bi-temporal or short-sequence comparisons. Moreover, current multi-temporal MLLMs in remote sensing generally operate at a high semantic level, lacking the pixel-precise analysis capabilities crucial for quantitative change assessment. These limitations pose significant challenges for dynamic urban understanding applications, which require both long time-series understanding beyond bi-temporal inputs and precise quantitative analysis of environmental changes.
@@
To address these challenges, we present DVL-Suite, a comprehensive framework for advancing long-term urban dynamics analysis through remote sensing imagery. The framework first introduces DVL-Bench, a large-scale benchmark for systematic evaluation. Inspired by recent advances in video understanding benchmarks ren2024timechat; liu2024bench; yuan2025videorefersuite, we establish a systematic taxonomy comprising seven essential capabilities for sustainable urban understanding (Fig. 1): 1) [BCA] Basic Change Analysis: Focuses on identifying and comparing multi-temporal changes in land-use patterns. 2) [CSE] Change Speed Estimation: Tracks and quantifies temporal trends of key urban elements, such as building expansion rates and vegetation loss. 3) [EA] Environmental Assessment: Evaluates urban livability and economic indicators through visual analysis. 4) [RCD] Referring Change Detection: Tests models’ capabilities in dense reasoning and precise spatial localization of changes. 5) [RCC] Regional Changes Caption: Generates detailed change descriptions for user-specified geographical areas. 5) [DTC] Dense Temporal Caption: Generates comprehensive reports documenting long-term temporal changes, highlighting critical events across long time series.
@@
To ensure comprehensive coverage, our benchmark encompasses diverse urban scenarios, including urban expansion, housing crises, natural disasters, urban heat island effects, and green space development. Unlike traditional bi-temporal understanding, our benchmark enables systematic evaluation of long time-series city dynamics, facilitating deeper insights into sustainable city development through multi-temporal Earth observation.
@@
Through extensive experimentation on DVL-Bench, we discovered that state-of-the-art MLLMs, both commercial and open-source models, face significant challenges in long-term temporal visual understanding, primarily due to insufficient training data spanning extended time periods. To address these limitations, we introduce DVL-Instruct, a specialized instruction-tuning dataset designed for dynamic city understanding in remote sensing. Building upon this dataset, we develop DVLChat, a baseline model that advances the capabilities of multi-temporal city understanding, captioning generation, and referring change detection.
@@
    We introduce DVL-Suite, comprising DVL-Bench and DVL-Instruct, with 15,063 high-resolution (1.0m) images spanning 42 U.S. cities, featuring an average of 6.85-6.94 temporal frames per scene from 2005 to 2023, enabling long-term urban dynamics analysis with a coherent task taxonomy, multi-level analysis capabilities, and thematic focus on urban development patterns.
@@
    We evaluate 17 vision-language models, revealing critical limitations with the best model, o4-mini, achieving only 34.1% accuracy on multiple-choice tasks on DVL-Bench, showing a significant gap in complex temporal tasks and quantitative analysis.
@@ 2.1 Large Multimodal Language Models
The rapid advancement of Multimodal Large Language Models (MLLMs) has sparked significant interest in their applications to complex visual understanding tasks, including understanding temporal dynamics and multiple image inputs, which are central to multi-temporal remote sensing analysis. In the field of generic vision-language models, early pioneering works such as Flamingo alayrac2022flamingo demonstrated the capability to process interleaved sequences of visual and textual data, including video frames, through mechanisms like Perceiver Resampler. Subsequent developments have focused on enhancing video understanding, with models like Video-LLaVA lin2024video unifying image and video representations before LLM projection. LLaVA-OneVision li2024llava-ov offers a unified framework for single-image, multi-image, and video via the proposed "Higher AnyRes" strategy. Qwen2-VL series wang2024qwen2vl; bai2025qwen25vl introduced Multimodal Rotary Position Embedding (M-ROPE) aligned with absolute time for long video comprehension, and InternVL3 zhu2025internvl3 employed Variable Visual Position Encoding (V2PE) for extended multimodal contexts including lengthy video sequences. Concurrently, the challenge of multi-image understanding has been addressed by models such as LLaVA-NeXT-Interleave li2025llavanextinterleave, which utilizes a data-centric approach with the M4-Instruct dataset to handle diverse multi-image scenarios. While existing MLLMs have progressed in handling temporal and multi-image data, they still fall short in dynamic, long-time-series remote sensing analysis, especially for precise quantitative assessment of urban changes. Therefore, we introduce the DVL-Suite and DVLChat to push the boundary of multi-temporal dynamic city understanding.
@@
Dataset Self- contained Average Temp. Text Pairs MT Images Image Size [BSC] [CRE] [EA] [RS] [RCC] [DTC]
@@
EarthDial soni2024earthdial $\times$ 1.01 11M 64.6k 448$\sim$1024 $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
DVL-Bench 6.85 8609 3469 1024 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
DVL-Instruct 6.94 61317 11594 1024 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@ 2.2 Multimodal Benchmarks in Remote Sensing
Remote Sensing (RS) domain has witnessed the emergence of numerous specialized vision-language datasets. LHRS-Bot muhtar2024lhrs, VRSBench livrsbench, and GeoChatSet kuckreja2024geochat pioneered single-temporal instruction datasets for classification, detection, and visual question answering (VQA). Subsequently, CDVQA yuan2022change introduced change-aware VQA, while LEVIR-MCI Liu_Change_Agent integrated pixel-level masks. GeoLLaVA elgendy2024geollava and CC-Expert wang2024ccexpert enhanced interactive bi-temporal captioning, and OVG-360k li2024show provided fine-grained spatial semantic supervision. Although surpassing single-temporal analyses, these efforts remain limited to bi-temporal image pairs. Recently, TeoChatlas irvin2024teochat has explored multi-temporal tasks derived from xView gupta2019xbd and fMoW christie2018functional. However, datasets with long time series remain sparse, and existing tasks primarily focus on bi-temporal classification using limited-resolution imagery. To enable MLLM to excel in understanding long time-series remote sensing images, we introduce DVL-Bench, a large-scale vision-language benchmark for remote sensing analysis within long time series that offers three key advantages: 1) Coherent task taxonomy. Unlike composite datasets that repurpose existing diverse labeled data, DVL-Bench features purpose-built data collection with a comprehensive and taxonomic coherence QA framework. 2) Diverse temporal tasks. DVL-Bench includes multiple analysis levels, from pixel-precise change tracking and regional dynamic captioning to comprehensive temporal understanding and environmental assessment, enabling systematic urban understanding. 3) Practical and thematically focused. In contrast to existing datasets addressing wide-ranging geospatial tasks, DVL-Bench specifically targets the analysis and representation of long-term urban development dynamics. The developed DVLChat can be directly integrated with the NAIP platform, serving as AI assistants for urban understanding applications.
3 Data Curation Pipeline
@@
To ensure data diversity and quality, we built our dataset using high-resolution (1m GSD) remote sensing imagery from the National Agriculture Imagery Program (NAIP), covering 42 major U.S. cities. The imagery was first geo-referenced and processed into 3907 patches of $1024\times 1024$ pixels, comprising 573 multi-temporal scenes. For compatibility with generic MLLMs, we utilized three optical bands, excluding the near-infrared band. By collecting environmental datasets from diverse Earth observation platforms (sources are provided in Appendix §D), all data were spatially resampled to match the resolution of our remote sensing imagery, ensuring one-to-one correspondence between images and environmental indicators. Based on the multi-source Earth observation data, we hired several experts and a well-trained annotation team to label and examine the DVL-Bench.
@@
Figure 2: The annotation pipeline of DVL-Bench dataset. Four common urban dynamics are depicted from the top to bottom, i.e., partial urban reconstruction, natural disaster, farmland conversion, and homeless encampments. In our semi-auto pipeline, the urban experts perform the basic annotations, and GPT4.1 integrates information and generates the paired instructions.
@@
Figure 2 illustrates our multi-stage annotation pipeline for the DVL-Bench dataset. Our annotation process encompasses several specialized tasks: For bi-temporal analysis tasks ([BCA] and [CSE]), annotators first segmented semantic change areas between adjacent temporal images. Changes were categorized across five primary land-cover types (vegetation, non-vegetated surfaces, water, buildings, and playgrounds), resulting in 20 distinct change event categories. Using these segmentation masks, GPT-4.1 generated diverse task-specific instructions. For [BCA] questions (e.g., "What land-cover type changed most between 2015 and 2017?"), the system calculated correct answers from the masks and generated four incorrect options from other land-cover types. For [CSE] tasks, change speeds were computed from the masks, with alternative options varying by ±20% and ±40%. The [EA] task followed a similar pipeline, but utilized our multi-source environmental indicators as reference data. For [RCD] tasks, domain experts designed event-specific prompts, followed by manual mask annotation and GPT-4.1-based language enhancement of prompts and answers. For temporal narrative tasks ([DTC] and [RCC]), annotators first identified keyframes containing significant changes to segment the temporal sequence, then crafted period-specific captions. GPT-4.1 refined these draft captions to enhance detail and coherence. While both tasks follow identical procedures, [RCC] focuses on user-specified local areas rather than global changes.
@@
4 DynamicVL Benchmark
DVL-Bench consists of 3,469 multi-temporal Earth observation images, each accompanied by human-verified annotations. The annotation suite includes 1,318 referring segmentation instructions for precise change localization, 5,854 question-answer pairs for detailed temporal analysis, and 1,437 comprehensive captions documenting urban dynamics. This section outlines our task taxonomy and examines the fundamental challenges in interpreting long-term remote sensing sequences.
Instruction distributions with different tasks. Fig. 4 presents the distribution of instructions across task categories. The [BCA], [CSE], [EA], and [RCD] tasks exhibit relatively balanced sample distributions, collectively accounting for 89.9% of the benchmark. In contrast to these tasks, which primarily focus on part of frames within each scene, the [DTC] and [RCC] tasks require comprehensive analysis of complete temporal sequences, thus representing smaller proportions of the overall distribution. With substantial sample sizes across all categories, our benchmark enables systematic evaluation of MLLMs in both dynamic understanding and generation performances.
Geospatial and temporal distributions. As shown in Fig. 4, the spatial and temporal distribution of our dataset. The samples are evenly distributed in 42 rapidly growing cities, ensuring comprehensive geographical coverage and minimizing regional bias. Unlike the existing multi-temporal dataset focusing on understanding bi-temporal changes, the DVL-Bench features long time-series analysis with sequences ranging from five to ten frames per urban scene. This extended temporal scope poses new challenges for modeling discrete temporal transitions in urban environments.
Figure 3: Task taxonomy and sample distribution in DVL-Bench. Multi-level task evaluates MLLM comprehensively.

### v2

# Weihao Xuan1,2∗  Junjue Wang1    Heli Qi2,3  Zihang Chen4 Zhuo Zheng5Yanfei Zhong4Junshi Xia2Naoto Yokoya1,2

@@ Abstract
Multimodal large language models (MLLMs) have demonstrated remarkable capabilities in visual understanding, but their application to long-term Earth observation analysis remains limited, primarily focusing on single-temporal or bi-temporal imagery. To address this gap, we introduce DVL-Suite, a comprehensive framework for analyzing long-term urban dynamics through remote sensing imagery. Our suite comprises 14,871 high-resolution (1.0m) multi-temporal images spanning 42 major cities in the U.S. from 2005 to 2023, organized into two components: DVL-Bench and DVL-Instruct. The DVL-Bench includes six urban understanding tasks, from fundamental change detection (pixel-level) to quantitative analyses (regional-level) and comprehensive urban narratives (scene-level), capturing diverse urban dynamics including expansion/transformation patterns, disaster assessment, and environmental challenges. We evaluate 18 state-of-the-art MLLMs and reveal their limitations in long-term temporal understanding and quantitative analysis. These challenges motivate the creation of DVL-Instruct, a specialized instruction-tuning dataset designed to enhance models’ capabilities in multi-temporal Earth observation. Building upon this dataset, we develop DVLChat, a baseline model capable of both image-level question-answering and pixel-level segmentation, facilitating a comprehensive understanding of city dynamics through language interactions. Project: https://github.com/weihao1115/dynamicvl.
@@ 1 Introduction
Sustainable city, as a key goal in “The 2030 Agenda for Sustainable Development” [^1] https://sdgs.un.org/goals/goal11, has proposed new requirements for urban resilience, convenience, and comfort. Remote sensing technology enables us to monitor urban development over time by analyzing satellite imagery, allowing us to track large-scale changes in urban landscapes frolking2024global; xiong2024earthnets. However, research in this field has been largely limited to comparing images from only two time points zhengsegment; wu2023fully, primarily due to the scarcity of well-aligned vision-language datasets spanning longer time series. This limitation has constrained our ability to conduct a comprehensive, large-scale understanding of urban dynamics.
@@
The recent emergence of MLLMs li2024llava-ov; liu2024llavanext; wang2024qwen2vl represents a significant advancement in visual-language understanding. These models mark a shift from specialized, single-purpose systems to versatile frameworks capable of handling multiple tasks, including but not limited to visual grounding liu2024bench, image captioning dai2023instructblip, and visual question answering yue2024mmmu; yue2024mmmupro. While recent MLLM research has demonstrated promising results in multi-image meng2024mmiu; jiangmantis and video understanding wu2024longvideobench; hu2025videommmu, these efforts primarily focus on in-the-wild daily imagery. In the context of remote sensing, existing research irvin2024teochat; soni2024earthdial on multi-temporal analysis is typically limited to bi-temporal or short-sequence comparisons. Moreover, current multi-temporal MLLMs in remote sensing are primarily tested on high-level semantic understanding, lacking pixel-precise analysis capabilities crucial for quantitative change assessment. These limitations pose significant challenges for dynamic city understanding applications, which require both long-term understanding beyond bi-temporal inputs and precise quantitative analysis of environmental changes.
@@
To address these challenges, we present DVL-Suite, a comprehensive framework for analyzing urban dynamics over time using remote sensing imagery. Our framework introduces DVL-Bench, a large-scale benchmark designed for rigorous evaluation of vision-language models in urban contexts. Building on recent advances in video understanding benchmarks ren2024timechat; liu2024bench; yuan2025videorefersuite, we develop a structured taxonomy that identifies six core capabilities essential for sustainable urban understanding: 1) [BCA] Basic Change Analysis: Focuses on identifying and comparing multi-temporal changes in land-use patterns. 2) [CSE] Change Speed Estimation: Tracks and quantifies temporal trends of key urban elements, such as building expansion rates and vegetation loss. 3) [EA] Environmental Assessment: Evaluates urban livability and economic indicators through visual analysis. 4) [RCD] Referring Change Detection: Tests models’ capabilities in dense reasoning and precise spatial localization of changes. 5) [RCC] Regional Change Captioning: Generates detailed change descriptions for user-specified geographical areas. 6) [DTC] Dense Temporal Captioning: Generates comprehensive reports documenting long-term temporal changes, highlighting critical events across long time series.
@@
To ensure comprehensive coverage, DVL-Bench encompasses diverse urban scenarios, such as urban expansion, housing crises, natural disasters, urban heat island effects, and green space development. Unlike traditional bi-temporal understanding, it enables systematic evaluation of long-term city dynamics, facilitating deeper insights into sustainable city development through multi-temporal Earth observation.
@@
Through extensive experimentation on DVL-Bench, we discovered that state-of-the-art MLLMs, both commercial and open-source models, face significant challenges in long-term temporal visual understanding, primarily due to insufficient training data spanning extended time periods. To address these limitations, we introduce DVL-Instruct, a specialized instruction-tuning dataset designed for dynamic city understanding in remote sensing. Using this dataset, we develop DVLChat, a baseline model that enhances multi-temporal urban understanding and caption generation, while introducing referring change detection capabilities.
@@
    We introduce DVL-Suite, comprising DVL-Bench and DVL-Instruct, with 14,871 high-resolution (1.0m) images spanning 42 U.S. cities, featuring an average of 6.73-6.94 temporal frames per scene from 2005 to 2023, enabling long-term urban dynamics analysis with a coherent task taxonomy, multi-level analysis capabilities, and thematic focus on urban development patterns.
@@
    We evaluate 18 vision-language models, revealing critical limitations: the best-performing model, o4-mini, achieves only 34.1% accuracy on DVL-Bench’s overall QA average, demonstrating significant deficiencies in complex temporal tasks and quantitative analysis.
@@ 2.1 Large Multimodal Language Models
The rapid advancement of MLLMs has sparked significant interest in their applications to complex visual understanding tasks, including understanding temporal dynamics and multiple image inputs, which are central to multi-temporal remote sensing analysis. In generic vision-language models, early pioneering works such as Flamingo alayrac2022flamingo demonstrated the capability to process interleaved sequences of visual and textual data, including video frames, through mechanisms like Perceiver Resampler. Subsequent developments have enhanced video understanding, with models like Video-LLaVA lin2024video unifying image and video representations before LLM projection, and LLaVA-OneVision li2024llava-ov offering a unified framework for single-image, multi-image, and video tasks via the "Higher AnyRes" strategy. Qwen2-VL series wang2024qwen2vl; bai2025qwen25 introduced Multimodal Rotary Position Embedding (M-ROPE) aligned with absolute time for long video comprehension, and InternVL3 zhu2025internvl3 employed Variable Visual Position Encoding (V2PE) for extended multimodal contexts including lengthy video sequences. Concurrently, the challenge of multi-image understanding has been addressed by models such as LLaVA-NeXT-Interleave li2025llavanextinterleave, which utilizes a data-centric approach with the M4-Instruct dataset to handle diverse multi-image scenarios.
Despite these advances in temporal and multi-image processing, existing models still fall short in dynamic, long-term remote sensing analysis, particularly for precise quantitative assessment of urban changes. To address this gap, we introduce DVL-Suite and DVLChat to advance multi-temporal urban understanding.
@@
Dataset
Self-
contained
Average
Temp.
Text
Pairs
MT
Images
Image
Size
[BCA] [CSE] [EA] [RCD] [RCC] [DTC]
@@
[…]
DVL-Bench 6.94 8,682 3,469 1024 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
DVL-Instruct 6.73 63,771 11,402 1024 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@ 2.2 Multimodal Benchmarks in Remote Sensing
Remote Sensing (RS) domain has witnessed the emergence of numerous specialized multimodal datasets xiao2025foundation. LHRS-Bot muhtar2024lhrs, VRSBench livrsbench, and GeoChatSet kuckreja2024geochat pioneered single-temporal instruction datasets for classification, detection, and visual question answering (VQA). Subsequently, CDVQA yuan2022change introduced change-aware VQA, while LEVIR-MCI Liu_Change_Agent integrated pixel-level masks. GeoLLaVA elgendy2024geollava and CC-Expert wang2024ccexpert enhanced interactive bi-temporal captioning, and OVG-360k li2024show provided fine-grained spatial semantic supervision. Although surpassing single-temporal analyses, these efforts remain limited to bi-temporal image pairs. Recently, TEOChatlas irvin2024teochat curates temporal instruction-following tasks such as those derived from xBD gupta2019xbd and fMoW christie2018functional. DisasterM3 wang2025disasterm3 provides a multi-hazard, multi-sensor, and multi-task remote sensing vision-language benchmark with 26,988 bi-temporal satellite images and diverse disaster assessment tasks. However, existing datasets predominantly focus on bi-temporal understanding and lack comprehensive evaluations of models’ capabilities in processing extended temporal sequences and performing long-term spatiotemporal reasoning. To enable MLLM to excel in understanding long-term remote sensing images, we introduce DVL-Bench, a large-scale vision-language benchmark for remote sensing analysis within long time series that offers three key advantages: 1) Coherent task taxonomy. Unlike composite datasets assembled from heterogeneous sources, DVL-Bench introduces a systematically designed task taxonomy built upon newly collected data with consistent annotation standards. 2) Diverse temporal tasks. DVL-Bench includes multiple analysis granularities, progressing from fine-grained pixel-level change detection and region-based dynamic captioning to holistic temporal reasoning and comprehensive environmental assessment, thereby facilitating systematic city dynamics understanding. 3) Practical and thematically focused. In contrast to existing datasets addressing wide-ranging geospatial tasks, DVL-Bench specifically targets the analysis and representation of long-term urban development dynamics. In addition, the developed DVLChat can be directly integrated with the NAIP platform, serving as an AI assistant for urban understanding applications.
3 DVL-Suite Curation Pipeline
@@
To ensure data diversity and quality, we built DVL-Suite using high-resolution (1.0m GSD) remote sensing imagery from the National Agriculture Imagery Program (NAIP), covering 42 major U.S. cities. The imagery was first geo-referenced and processed into 14,871 patches of $1024\times 1024$ pixels, comprising 2,193 multi-temporal scenes. For compatibility with generic MLLMs, we utilized three optical bands. By collecting environmental datasets from diverse Earth observation platforms (sources are detailed in Appendix § D), all data were spatially resampled to match the resolution of collected remote sensing imagery, ensuring one-to-one correspondence between images and environmental indicators. Based on the multi-source Earth observation data, we hired several experts and a well-trained annotation team to label and examine the DVL-Suite.
@@
Figure 2: The annotation pipeline of the proposed DVL-Suite. Four common urban dynamics are depicted from top to bottom: partial urban reconstruction, natural disasters, farmland conversion, and homeless encampments. In our semi-auto pipeline, urban experts perform the basic annotations, while GPT4.1 integrates this information to generate the paired instructions.
@@
Figure 2 illustrates our multi-stage annotation pipeline for the DVL-Bench dataset. The annotation process includes several specialized tasks: For bi-temporal analysis tasks ([BCA] and [CSE]), annotators first segmented semantic change areas between adjacent temporal images. These changes were categorized across five primary land-cover types: vegetation, non-vegetated surfaces, water, buildings, and playgrounds. This categorization, adapted from SECOND yang2020semantic, yielded 20 distinct change event categories. GPT4.1 then generated diverse task-specific instructions using these segmentation masks and categories. For [BCA] questions (e.g., "What land-cover type changed most between 2015 and 2017?"), the system calculated correct answers from the masks and generated four incorrect options from other land-cover types. For [CSE] tasks, change speeds were computed from the masks, with alternative options varying by ±20% and ±40%. The [EA] task followed a similar pipeline, but utilized the multi-source environmental indicators as reference data. For [RCD] tasks, domain experts designed event-specific prompts, followed by manual mask annotation and GPT4.1-based language enhancement of prompts and answers. For temporal narrative tasks ([DTC] and [RCC]), annotators first identified keyframes containing significant changes to segment the temporal sequence, then crafted period-specific captions. GPT4.1 refined these draft captions to enhance detail and coherence. While both tasks follow identical procedures, [RCC] focuses on user-specified local areas rather than global changes.
@@
4 DVL-Bench Designs
DVL-Bench consists of 3,469 multi-temporal Earth observation images, each accompanied by human-verified annotations. The annotation suite includes 1,391 referring segmentation instructions for precise change localization, 5,854 question-answer pairs for detailed temporal analysis, and 1,437 comprehensive captions documenting urban dynamics. This section outlines the task taxonomy and examines the fundamental challenges in interpreting long-term remote sensing sequences.
Instruction distributions with different tasks. Figure 4 presents the distribution of instructions across task categories. The [BCA], [CSE], [EA], and [RCD] tasks exhibit relatively balanced sample distributions, collectively accounting for 89.9% of the benchmark. In contrast to these tasks, which primarily focus on part of frames within each scene, the [DTC] and [RCC] tasks require comprehensive analysis of complete temporal sequences, thus representing smaller proportions of the overall distribution. With substantial sample sizes across all categories, DVL-Bench enables systematic evaluation of MLLMs in both dynamic understanding and open-ended generation performances.
Geospatial and temporal distributions. Figure 4 shows the spatial distribution across 42 U.S. cities and the temporal length per scene. The samples are evenly distributed in 42 rapidly growing cities, ensuring comprehensive geographical coverage and minimizing regional bias. Unlike the existing multi-temporal dataset focusing on understanding bi-temporal changes, the DVL-Bench features long-term analysis with sequences ranging from five to ten frames per urban scene. This extended temporal scope poses new challenges for modeling discrete temporal transitions in urban environments.
Figure 3: Task taxonomy and sample distribution in DVL-Bench. The multi-level task evaluates MLLM comprehensively.